# Graph Neural Networks Gone Hogwild

### v1

@@ 3.2 Optimization-based GNNs
Optimization-based GNNs are an important special case of fixed-point GNNs, obtaining node embeddings by minimizing a convex scalar-valued graph function ${E_{\theta}:{\mathcal{G}}\times\mathbb{R}^{n\times k}\to\mathbb{R}}$ with respect to the aggregate graph node embeddings ${\bm{H}}$, where $\theta$ are parameters:
@@
In this work, we assume $E_{\theta}$ is separable per node, with each node computing its contribution to $E_{\theta}$ using local information. That is, $E_{\theta}$ can be expressed as:
@@
Here $e^{i}_{\theta}$ depends only on information from node $i$’s neighborhood, and we further assume it is twice differentiable with a Hessian of bounded norm. Crucially, due to the dependence of each $e^{i}_{\theta}$ on local information, gradient-based minimization of $E_{\theta}$ can be expressed per node via message passing:
@@
where ${g_{\theta}:\mathbb{R}^{n\times p}\rightarrow\mathbb{R}^{n\times k}}$, ${\bm{L}}$ is a function of ${\bm{A}}$ and can be viewed as a generalized incidence matrix (assumed to be symmetric and positive semi-definite), $\gamma$ and $\beta$ are constants, and $\text{tr}()$ is the trace. The first term drives node embeddings to approximate some function of the node features ${\bm{X}}$, and the second term is a regularizer that rewards smoothness of neighboring node embeddings. We refer to optimization-based GNNs which use this form of objective as GSDGNNs, since minimizing the objective can be interpreted as performing graph signal denoising (GSD).
Previous work has shown that the node embeddings obtained by the fixed-point message passing scheme defined by APPNP (Gasteiger et al., 2019) and by EIGNN (Liu et al., 2021) correspond directly to minimization of this form of objective (Ma et al., 2021; Yang et al., 2021). This correspondence between fixed-point GNNs and optimization-based GNNs does not always exist; while the node embeddings in optimization-based GNNs can always be expressed as satisfying a fixed-point equation (i.e., the gradient of the convex graph function is equal to zero at the solution), not all fixed-point equations correspond with convex optimization problems. For this reason, we distinguish between these two types of implicit GNNs.
4 Hogwild GNN Inference
Executing the layer updates during GNN inference is typically performed synchronously, as depicted in Figure 1; each node embedding is derived from the previous layer embeddings. In this section, we discuss GNN inference under the asynchronous, “hogwild” (Niu et al., 2011) execution model presented in Section 2.1.
In the general asynchronous update in Equation 2 it is assumed that node $i$ has access to all of the values ${\bm{h}}_{1},...{\bm{h}}_{n}$ required for performing its update. In general these values could be obtained, for example, by providing nodes access to a shared memory structure which contains these values. However, in this work, we are interested in decentralized, distributed, and asynchronous GNN embedding updates, where neither centralized memory nor a centralized controller are used. We assume that individual nodes store their own values ${\bm{h}}_{i}$ (and ${\bm{x}}_{j}$, ${\bm{e}}_{ij}$), and if these values are needed by other nodes to perform their update, they must obtain them by communicating with neighbors. Our aim in this section is to demonstrate that node updates can be performed only using information that can be obtained through local communication.
@@ 4.1 Explicitly-defined GNN and Fixed-point GNN Inference under Partial Asynchronism
Without loss of generality, assume that the embedding dimension $k$ is fixed for all layers of parameterized node update functions. We do not write $f_{\theta}$ indexed by layer, but this is straightforwardly generalized to the case of layer-specific parameters and functions described in Section 2.2. Without loss of generality, assume the embedding update function $f_{\theta}$ is continuously differentiable, so that the following restriction can be stated:
$j\notin\mathcal{N}(i)\;\implies\frac{\partial f_{\theta}}{\partial{\bm{h}}_{j}}({\bm{z}})=0\quad\forall{\bm{z}}\in\mathbb{R}^{k}.$ (8)
@@
where we omit node and edge features for clarity. Note the crucial difference introduced by asynchrony: the neighbor data ${\bm{h}}_{j}(\tau_{j}^{i}(t))$ may be out of date, corresponding to a different layer in the network. For explicitly-defined GNNs such as GCN or GAT, the number of iterations $|T_{i}|$ executed by each node is fixed and equal to the number of layers $L$ in the GNN. For fixed point GNNs, the number of iterations is not pre-specified.
Since explicitly-defined GNNs implement a specific feed-forward neural network architecture, partially asynchronous inference using Equation 9 corrupts the computation performed by the network. We illustrate this in Figure 1, where partial asynchrony results in a (different) computation graph with some connections removed, and new connections that are not present in the original synchronous computation graph. This means that there are no convergence guarantees under partial asynchrony, and in general the final node embeddings may vary significantly with respect to the particular node update sequence.
In contrast, fixed-point GNN architectures in which $F_{\theta}$ is contractive with respect to node embeddings ${\bm{h}}$ are provably robust to partially asynchronous inference. In particular, they satisfy the assumptions of the following proposition (Bertsekas, 1983).
@@ 4.2 Optimization-based GNN Inference under Partial Asynchrony
However, recall that for a node to perform an update using only local communication, we previously assumed that ${\bm{g}}_{ji}$ was obtained by node $i$ from its neighbor $j$. With the formulation in Equations 11 and 12, node $j$ cannot provide node $i$ with ${\bm{g}}_{ji}$ since the value depends on node $i$’s view of the embeddings (and features) of the neighbors of node $j$, rather than deferring to node $j$’s view of its neighbors. That is, node $i$ needs access to information about its 2-hop neighbors in addition to its 1-hop neighbors. Since we assume communication is only possible with 1-hop neighbors, 2-hop neighbor information would need to be forwarded by direct neighbors of node $i$. This inflates the cost of communication, requiring a number of bits per transmission which is proportional to $|\mathcal{N}(j)|$ Furthermore, as the number of neighbors that are shared between node $i$ and node $j$ increases, the contents of these messages become increasingly redundant.
@@
If $E_{\theta}$ is strongly convex and separable per node, then for a sufficiently small step size and the bounded staleness conditions in 2.1, the optimization procedure of Equations 11 and 13 will converge when executed in a distributed manner underp partial asynchrony.
@@ 5 Energy GNNs
Under the assumptions of Section 2.1, implicit GNNs in which node embeddings are updated iteratively using local information are well suited for partially asynchronous, decentralized, and distributed inference. However, as we discussed in Section 3, the diversity of existing implicit GNN architectures (both fixed-point and optimization-based) is limited. Fixed-point GNNs employ simple message-passing schemes that can be enforced to be contractive, where $f_{\theta}$ consists of a linear function to compute messages ${\bm{m}}_{i}$, possibly followed by the application of a component-wise non-decreasing non-linearity to compute updated embeddings. Optimization-based GNNs use a restricted class of graph objectives $E_{\theta}$ as shown in Equation 7. One limitation with this objective is that it assumes informative node features exist, which may not be the case. Edge features are not easily incorporated, which may be present and relevant for prediction. Finally, smoothness across embeddings for neighboring nodes may be inappropriate for some datasets or tasks where neighboring nodes are expected to generate distinct embeddings.
@@
$e^{i}_{\theta}$ $=u\left({\bm{m}}_{i},{\bm{h}}_{i},{\bm{x}}_{i};\theta_{u}\right)+(\nicefrac{{\beta}}{{2}})||{\bm{h}}_{i}||_{2}^{2}\,.$ (15)
@@
Here ${\bm{m}}_{i}$ is the aggregation of state from neighboring nodes, where $m$ is a function that is both convex and nondecreasing (in each dimension) with respect to ${\bm{h}}_{j}$ and ${\bm{h}}_{i}$. The function $u$ is convex with respect to ${\bm{m}}_{i}$ and ${\bm{h}}_{i}$. These functions are both implemented as PICNNs and their composition is convex. Summing these functions along with the squared norm penalty results in $E_{\theta}$ being strongly convex with respect to the node embeddings ${\bm{H}}$. This architecture for the energy can be described as a (single layer) partially input-convex GNN; we provide more details in Appendix B.
This formulation for $E_{\theta}$ offers significantly more flexibility than the architectures of other implicitly-defined GNNs. Any combination of inputs to $m$ or $u$ are valid, provided that $E_{\theta}$ remains a convex function of ${\bm{H}}$. This means that edge features are easily incorporated into the model, and neighbor-specific or neighbor agnostic messages can be used (e.g., $m$ can take in information from just a node’s neighbor, or information pertaining to both a node and its neighbor); in our experiments in Section 6.3 we show that this translates empirically to improved performance on various tasks. Since the aggregation in Equation 16 is only constrained to be a non-negative sum over neighbors, it can be replaced with, for example, a mean or a sum weighted by the entries of the symmetric renormalized adjacency matrix (as in GCNs, see Appendix A). Alternatively, Equation 16 can easily incorporate a neighbor attention mechanism (as in GATs), where neighbor contributions to the sum are scaled by neighbor-specific attention weights (see Appendix B). The attention weights can depend on any of the non-convex inputs (i.e., the features).
@@ 6.1 Synthetic Multi-Agent Tasks
We perform experiments on a number of synthetic datasets, motivated by tasks which are of interest for multi-agent systems and where distributed, asynchronous inference is desirable. We describe each task and associated dataset below.
@@
The ability to communicate information across long distances in a group of agents is important when agent predictions depend on global information. This communication is made more difficult in the absence of a central controller (as is the case for distributed, asynchronous inference).
The chains dataset, used in Gu et al. (2020); Liu et al. (2021), is meant to evaluate the ability to capture long-range dependencies between nodes. The dataset consists of $p$ undirected linear graphs with $l$ nodes, with each graph having a label $k\in\{1,...,p\}$. The task is node classification of the graph label, where class information is contained only in the feature of the first node in the chain; the node feature matrix ${\bm{X}}\in\mathbb{R}^{n\times p}$ for a graph with class $k$ has ${\bm{X}}_{1,k}=1$ and zeros at all other indices. Perfect classification accuracy indicates that information is successfully propagated from the first node to the final node in the chain. For our dataset, we use chains with length $l=100$.
@@
For this task, we consider summation, a basic functional building block relevant for many multi-agent tasks. For instance, in reinforcement learning tasks, agents might aim to perform actions that optimize their collective rather than individual rewards, requiring each agent to sum the rewards associated with all other agents. Many distributed and asynchronous algorithms exist for summation (Kempe et al., 2003).
We construct a dataset to evaluate the ability of GNNs to perform binary sums in undirected chain graphs. Our data are $2000$ graphs with $50$ nodes each, with different instantiations of binary node features ${\bm{x}}_{i}\in\{0,1\}$. The prediction target for each node in a given graph is $\hat{{\bm{y}}}_{i}:=\sum_{i}{\bm{x}}_{i}$.
@@
A common task for multi-agent collectives such as robot swarms is localization. This problem has previously been tackled in various ways that all employ a bespoke algorithm tailored for the task (Todescato et al., 2016; Huang and Tian, 2017, 2018).
We test the ability of GNNs to solve this problem on static graphs. We construct a dataset where each node has a position in $\mathbb{R}^{2}$ and neighbors within some radius are connected by an edge. We do not assume a global coordinate system; instead, we focus on relative localization, where pairwise distances between nodes are maintained. Each node predicts a position in $\mathbb{R}^{2}$, and the objective is the mean squared error between true pairwise node distances, and distances between their predicted positions. In order to break symmetries, each node has a unique ID which is one-hot encoded and used as the node feature. Distances to connected neighbors are provided as edge features. We generate $1500$ random graphs where all graphs consist of $20$ nodes. We sample uniformly in the unit square to get node positions and connect nodes by an edge if they are within a distance of 0.5.
@@
The final synthetic task we consider is “terrain” classification. Suppose a number of agents are placed in an environment where each agent performs some local measurement, and the agents must collectively make predictions about some global state of the environment using only local communication. For this experiment, we use MNIST images (those with 0/1 labels only) to represent the environment (LeCun et al., 2010). Agents are placed at random locations in the image, and use the coordinates and pixel value at their location as their node features. The prediction target for each agent is the image label. We resize the images to $10\times 10$ pixels, and sample 10 random pixels for agent locations. Nodes share an edge if they are within 5 pixels of each other.
Unlike the previous synthetic tasks, in which existing distributed algorithms can be applied, no bespoke algorithm exists for the MNIST terrain task. This is precisely the type of problem which motivates the development of GNNs which are robust to asynchronous and distributed inference.
@@ 6.2 Experimental Setup
The cost of the forward and backward pass (in terms of computation and/or memory) for implicitly-defined GNNs is variable, depending on the number of iterations required for convergence. We mitigate this cost during training in two ways. Since convergence for both fixed-point GNNs and optimization-based GNNs is guaranteed (since for fixed-point GNNs, $F_{\theta}$ is contractive and for optimization-based GNNs, $E_{\theta}$ is convex) implicit differentiation can be used to obtain gradients of the task-specific loss function $\mathcal{L}$ with respect to parameters $\theta$. This avoids unrolling the fixed-point iterations in the backward pass, and requires a fixed amount of computation and memory. We derive the gradient in Appendix C. Furthermore, since the solution of the forward pass is unique (i.e. not dependent on the initialization of ${\bm{H}}$), the number of iterations in the forward pass can be reduced by initializing ${\bm{H}}$ to be the solution from the previous epoch of training. In our work we employ both of these strategies during training.
@@
Additional training details for the synthetic experiments are in Appendix F.

### v2

@@ 3.2 Optimization-based GNNs
Optimization-based GNNs obtain node embeddings by minimizing a convex scalar-valued graph function ${E_{\theta}:{\mathcal{G}}\times\mathbb{R}^{n\times k}\to\mathbb{R}}$ with respect to the aggregate graph node embeddings ${\bm{H}}$, where $\theta$ are parameters:
@@
Assuming $E_{\theta}$ is separable per node, $E_{\theta}$ can be expressed as:
@@
Crucially, due to the dependence of each $e^{i}_{\theta}$ on only local information, gradient-based minimization of $E_{\theta}$ can be expressed per node via message passing:
@@
where ${g_{\theta}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{k}}$ and is applied independently to each node feature, ${\bm{L}}$ is a function of ${\bm{A}}$ and can be viewed as a generalized incidence matrix (assumed to be symmetric and positive semi-definite), $\gamma$ and $\beta$ are constants, and $\text{tr}()$ is the trace. The first term drives node embeddings to approximate some function of the node features ${\bm{X}}$, and the second term is a regularizer that rewards smoothness of neighboring node embeddings. The overall objective $E_{\theta}$ is separable per node. We refer to optimization-based GNNs which use this form of objective as GSDGNNs, since minimizing the objective can be interpreted as performing graph signal denoising (GSD).
Previous work has shown that embeddings obtained by the fixed-point message passing scheme defined by APPNP (Gasteiger et al., 2019) and EIGNN (Liu et al., 2021) correspond to minimization of this form of objective (Ma et al., 2021; Yang et al., 2021). Correspondence between fixed-point GNNs and optimization-based GNNs does not always exist; while node embeddings in optimization-based GNNs can always be expressed as satisfying a fixed-point equation (i.e., the gradient of the convex graph function is equal to zero at the solution), not all fixed-point equations correspond with convex optimization problems. For this reason, we distinguish between these two types of implicit GNNs.
4 Asynchronous GNN Inference
In this section, we discuss GNN inference under the asynchronous execution model presented in Section 2.1. In the general asynchronous update in Equation 2 it is assumed that node $i$ has access to all of the values ${\bm{h}}_{1},...{\bm{h}}_{n}$ required for performing its update. In general these values could be obtained, for example, by providing nodes access to a shared memory structure which contains these values. However, in this work, we are interested the setting where neither centralized memory nor a centralized controller are used. We assume that individual nodes store their own values ${\bm{h}}_{i}$ (and ${\bm{x}}_{j}$, ${\bm{e}}_{ij}$), and if these values are needed by other nodes to perform their update, they must obtain them by communicating with neighbors. Our aim in this section is to demonstrate that node updates can be performed only using information that can be obtained through local communication.
@@ 4.1 Explicitly-defined GNN and Fixed-point GNN Inference under Partial Asynchronism
Without loss of generality, assume that the embedding dimension $k$ is fixed for all layers of parameterized node update functions. We do not write $f_{\theta}$ indexed by layer, but this is straightforwardly generalized to the case of layer-specific parameters and functions described in Section 2.2. As a slight abuse of notation, let $f_{\theta}^{i}$ denote $f_{\theta}$ applied to node $i$’s neighborhood. Without loss of generality, assume the embedding update function $f_{\theta}$ is continuously differentiable, so that the following restriction can be stated:
$j\notin\mathcal{N}(i)\;\implies\frac{\partial f^{i}_{\theta}}{\partial{\bm{h}}_{j}}({\bm{z}})=0\quad\forall{\bm{z}}\in\mathbb{R}^{k}.$ (9)
@@
where we omit node and edge features for clarity. Note the crucial difference introduced by asynchrony: the neighbor data ${\bm{h}}_{j}(\tau_{j}^{i}(t))$ may correspond to the incorrect layer in the network. For explicitly-defined GNNs such as GCN or GAT, the number of iterations $|T_{i}|$ executed by each node is fixed and equal to the number of layers $L$ in the GNN. For fixed point GNNs, the number of iterations is not pre-specified.
Since explicitly-defined GNNs implement a specific feed-forward neural network architecture, inference using Equation 10 corrupts the computation performed by the network. We illustrate this in Figure 1, where asynchrony results in a (different) computation graph with some connections removed, and new connections that are not present in the original synchronous computation graph. This means that there are no convergence guarantees under partial asynchrony, and in general the final node embeddings may vary significantly with respect to the particular node update sequence.
In contrast, fixed-point GNNs in which $F_{\theta}$ is contractive with respect to embeddings ${\bm{h}}$ are provably robust to partially asynchronous inference. In particular, they satisfy the assumptions of the following proposition (Bertsekas, 1983).
@@ 4.2 Optimization-based GNN Inference under Partial Asynchrony
However, recall that for a node to perform an update using only local communication, we previously assumed that ${\bm{g}}_{ji}$ was obtained by node $i$ from its neighbor $j$. With the formulation in Equations 12 and 13, node $j$ cannot provide node $i$ with ${\bm{g}}_{ji}$ since the value depends on node $i$’s view of the embeddings (and features) of the neighbors of node $j$, rather than deferring to node $j$’s view of its neighbors. That is, node $i$ needs access to information about its 2-hop neighbors in addition to its 1-hop neighbors. Since we assume communication is only possible with 1-hop neighbors, 2-hop neighbor information would need to be forwarded by direct neighbors of node $i$. This inflates the cost of communication, requiring a number of bits per transmission which is proportional to $|\mathcal{N}(j)|$. Furthermore, as the number of neighbors that are shared between node $i$ and node $j$ increases, the contents of these messages become increasingly redundant.
@@
If $E_{\theta}$ is strongly convex and separable per node, and is twice differentiable (w.r.t. ${\bm{H}}$) with a Hessian of bounded norm, then for a sufficiently small step size and the bounded staleness conditions in 2.1, the optimization procedure of Equations 12 and 14 will converge when executed under partial asynchrony.
@@ 5 Energy GNNs
Under the assumptions of Section 2.1, implicitly-defined GNNs in which node embeddings are updated iteratively using local information are well suited for partially asynchronous, decentralized, and distributed inference. However, the diversity of existing implicit GNN architectures is limited (see Section 3).
@@
$e^{i}_{\theta}$ $=u\left({\bm{m}}_{i},{\bm{h}}_{i},{\bm{x}}_{i};\theta_{u}\right)+(\nicefrac{{\beta}}{{2}})||{\bm{h}}_{i}||_{2}^{2}\,$ (16)
@@
where $m$ is a function that is both convex and nondecreasing (in each dimension) with respect to ${\bm{h}}_{j}$ and ${\bm{h}}_{i}$, and the function $u$ is convex with respect to ${\bm{m}}_{i}$ and ${\bm{h}}_{i}$. These functions are both implemented as PICNNs and their composition is convex. Summing these functions along with the squared norm penalty results in $E_{\theta}$ being strongly convex with respect to the node embeddings ${\bm{H}}$. This architecture for the energy can be described as a (single layer) partially input-convex GNN; we provide more details in Appendix B.
This formulation for $E_{\theta}$ offers significantly more flexibility than the architectures of other implicitly-defined GNNs. The functions $m$ and $u$ are parameterized by multi-layer PICNNs, and any combination of inputs to $m$ and $u$ are valid provided that $E_{\theta}$ remains a convex function of ${\bm{H}}$. This means that edge features are easily incorporated into the model, and neighbor-specific or neighbor agnostic messages can be used (e.g., $m$ can take in information from just a node’s neighbor, or information pertaining to both a node and its neighbor); in our experiments in Section 6.3 we show that this translates empirically to improved performance on various tasks. Since the aggregation in Equation 17 is only constrained to be a non-negative sum over neighbors, it can be replaced with, for example, a mean or a sum weighted by the entries of the symmetric renormalized adjacency matrix (as in GCNs, see Appendix A). Alternatively, Equation 17 can easily incorporate a neighbor attention mechanism (as in GATs), where neighbor contributions to the sum are scaled by neighbor-specific attention weights (see Appendix B). The attention weights can depend on any of the non-convex inputs (i.e., the features).
@@ 6.1 Synthetic Multi-Agent Tasks
We perform experiments on several synthetic datasets, motivated by prediction tasks which are of interest for multi-agent systems where distributed, asynchronous inference is desirable. We describe each task and associated dataset below.
@@
The ability to communicate information across long distances in a group of agents is important when agent predictions depend on global information. This communication is made more difficult in the absence of a central controller (as is the case for distributed, asynchronous inference). The chains dataset, used in Gu et al. (2020); Liu et al. (2021), is meant to evaluate the ability to capture long-range dependencies between nodes. The dataset consists of $p$ undirected linear graphs with $l$ nodes, with each graph having a label $k\in\{1,...,p\}$. The task is node classification of the graph label, where class information is contained only in the feature of the first node in the chain; the node feature matrix ${\bm{X}}\in\mathbb{R}^{n\times p}$ for a graph with class $k$ has ${\bm{X}}_{1,k}=1$ and zeros at all other indices. Perfect classification accuracy indicates that information is successfully propagated from the first node to the final node in the chain. For our dataset, we use chains with length $l=100$.
@@
For this task, we consider summation, a basic functional building block relevant for many multi-agent tasks. For instance, in reinforcement learning tasks, agents might aim to perform actions that optimize their collective rather than individual rewards, requiring each agent to sum the rewards associated with all other agents. Many distributed and asynchronous algorithms exist for summation (Kempe et al., 2003). We construct a dataset to evaluate the ability of GNNs to perform binary sums in undirected chain graphs. Our data are $2000$ graphs with $50$ nodes each, with different instantiations of binary node features ${\bm{x}}_{i}\in\{0,1\}$. The prediction target for each node in a given graph is $\hat{{\bm{y}}}_{i}:=\sum_{i}{\bm{x}}_{i}$.
@@
A common task for multi-agent collectives such as robot swarms is localization. This problem has previously been tackled in various ways that all employ a bespoke algorithm tailored for the task (Todescato et al., 2016; Huang and Tian, 2017; 2018). We test the ability of GNNs to solve this problem on static graphs. We construct a dataset where each node has a position in $\mathbb{R}^{2}$ and neighbors within some radius are connected by an edge. We do not assume a global coordinate system; instead, we focus on relative localization, where pairwise distances between nodes are maintained. Each node predicts a position in $\mathbb{R}^{2}$, and the objective is the mean squared error between true pairwise node distances, and distances between their predicted positions. In order to break symmetries, each node has a unique ID which is one-hot encoded and used as the node feature. Distances to connected neighbors are provided as edge features. We generate $1500$ random graphs where all graphs consist of $20$ nodes. We sample uniformly in the unit square to get node positions and connect nodes by an edge if they are within a distance of 0.5.
@@
The final synthetic task we consider is “terrain” classification. Suppose a number of agents are placed in an environment where each agent performs some local measurement, and the agents must collectively make predictions about some global state of the environment using only local communication. For this experiment, we use MNIST images (those with 0/1 labels only) to represent the environment (LeCun et al., 2010). Agents are placed at random locations in the image, and use the coordinates and pixel value at their location as their node features. The prediction target for each agent is the image label. We resize the images to $10\times 10$ pixels, and sample 10 random pixels for agent locations. Nodes share an edge if they are within 5 pixels of each other. Unlike the previous synthetic tasks, in which existing distributed algorithms can be applied, no bespoke algorithm exists for the MNIST terrain task. This is precisely the type of problem which motivates the development of GNNs which are robust to asynchronous and distributed inference.
@@ 6.2 Experimental Setup
The cost of the forward and backward pass (in terms of computation and/or memory) for implicitly-defined GNNs is variable, depending on the number of iterations required for convergence. We mitigate this cost during training in two ways. Since convergence for both fixed-point GNNs and optimization-based GNNs is guaranteed implicit differentiation can be used to obtain gradients of the task-specific loss function $\mathcal{L}$ with respect to parameters $\theta$. This avoids unrolling the fixed-point iterations in the backward pass, and requires a fixed amount of computation and memory. We derive the gradient in Appendix C. Furthermore, since the solution of the forward pass is unique (i.e. not dependent on the initialization of ${\bm{H}}$), the number of iterations in the forward pass can be reduced by initializing ${\bm{H}}$ to be the solution from the previous epoch of training. In our work we employ both of these strategies during training. Additional training details for the synthetic experiments are in Appendix F.